Would the answer be A?
9. An ammeter does not apprecby l curent of the circuit it is placed in because it is connected in series and has a low resistance y it is connected in parallel and has a high resistance it is connected in parallel and has a low resistance D) it is connected in series and has a high resistance it is connected in either series or parallel and has no resistance
Answer #1

Yes, the answer is A because If you are measuring any quantity in the circuit then there should not be any loss due to your instruments that you are using to measure something.

If ammeter has high resistance let's say r_{1} and the circuit in which you are measuring the current has resistance R then current in the circuit is

I_{1}=\frac{E}{R+r_{1}}

now if ammeter has resistance r_{2} that is small then current

I_{2}=\frac{E}{R+r_{2}}

since r_{2} is small we can neglect r_{2} so current

I_{2}\approx \frac{E}{R}

so I_{2}>I_{1}

That means for measuring current theammeter should have very small voltage so that there is no loss in the ammeter and it behaves like a short circuit.